### v1

# Urban4D: Semantic-Guided 4D Gaussian Splatting for Urban Scene Reconstruction

Abstract
Reconstructing dynamic urban scenes presents significant challenges due to their intrinsic geometric structures and spatiotemporal dynamics. Existing methods that attempt to model dynamic urban scenes without leveraging priors on potentially moving regions often produce suboptimal results. Meanwhile, approaches based on manual 3D annotations yield improved reconstruction quality but are impractical due to labor-intensive labeling. In this paper, we revisit the potential of 2D semantic maps for classifying dynamic and static Gaussians and integrating spatial and temporal dimensions for urban scene representation. We introduce Urban4D, a novel framework that employs a semantic-guided decomposition strategy inspired by advances in deep 2D semantic map generation. Our approach distinguishes potentially dynamic objects through reliable semantic Gaussians. To explicitly model dynamic objects, we propose an intuitive and effective 4D Gaussian splatting (4DGS) representation that aggregates temporal information through learnable time embeddings for each Gaussian, predicting their deformations at desired timestamps using a multilayer perceptron (MLP). For more accurate static reconstruction, we also design a k-nearest neighbor (KNN)-based consistency regularization to handle the ground surface due to its low-texture characteristic. Extensive experiments on real-world datasets demonstrate that Urban4D not only achieves comparable or better quality than previous state-of-the-art methods but also effectively captures dynamic objects while maintaining high visual fidelity for static elements.
Figure 1: Qualitative comparison on the nuScenes [4] dataset. While DeformGS [40] achieves comparable results on static regions, it fails on dynamic objects, producing severe artifacts and blurred reconstructions. In contrast, our Urban4D maintains high fidelity for both dynamic objects and static backgrounds, also surpassing the reconstruction quality of PVG [7].
Figure 2: Different from previous 3DGS-based methods which leverage pre-trained segmentation models to predict 2D semantic maps for downstream tasks, we revisit the potential of these estimated semantic maps to handle potentially dynamic objects for reconstruction. From left to right: input image, semantic map, our reconstruction, and reconstruction without potentially dynamic objects. Our method effectively preserves the geometry and appearance of both static structures and potentially dynamic objects.
1 Introduction
Recent advancements in urban scene reconstruction have witnessed remarkable progress, particularly with the introduction of 3D Gaussian Splatting (3DGS)  [13], which has enhanced real-time rendering while maintaining high reconstruction quality for static scenes. However, directly extending these capabilities to model dynamic and large-scale urban scenes presents significant challenges due to spatiotemporal dynamics and intricate geometric structures.
Urban scenes can generally be categorized into two types: dynamic objects, such as vehicles and pedestrians, which exhibit complex motion patterns, and static elements, which maintain consistent spatial relationships. Specifically, low-texture regions (e.g., ground surfaces) pose challenges for accurate static reconstruction due to their lack of salient patterns. These unique characteristics make reconstructing dynamic objects in urban scenes particularly difficult. The coexistence of static and dynamic elements necessitates a model that can distinguish between and process them independently to ensure accurate reconstruction.
Previous dynamic NeRF-based methods, such as  [38, 32], attempt to address these challenges by introducing time-conditioned neural networks to model motion, but they suffer from slow rendering speeds and limited capacity to handle complex and large-scale urban scenes. Some methods explore suitable representations, such as Periodic Vibration Gaussian (PVG)  [7], which proposes a unified framework by incorporating periodic vibration-based temporal dynamics in urban scenes. However, such a unified representation may degrade performance when treating static elements with temporal processes. Other 3DGS-based methods resort to manually labeled 3D bounding boxes to process dynamic objects  [37, 41], which is labor-intensive and impractical for large-scale or continuously changing environments.
Furthermore, we have witnessed that some of these methods adopt pre-trained segmentation models, such as SegFormer [35], for semantic novel view rendering. Semantic segmentation based on deep learning provides more reliable classifications of scene components in 2D images. Leveraging these results, we may effectively classify dynamic and static Gaussians without relying on manually labeled 3D annotations. The observations in Figure 2 lead us to an intriguing question: Can we revisit the potential of 2D semantic maps for classifying potentially dynamic objects and leverage spatiotemporal information to adequately represent urban scenes? The answer to this question uncovers an unexplored area in reconstructing dynamic objects based on 4D Gaussian splatting (4DGS), enabling a more holistic and precise representation.
Inspired by this perspective, we propose Urban4D, a novel framework to overcome the limitations of traditional 3DGS-based methods in handling temporal dynamics and reconstructing urban scenes. Specifically, we design a semantic-guided decomposition strategy, which builds on recent progress in deep 2D semantic map generation to distinguish potentially dynamic objects from static elements by assigning reliable semantic classes to Gaussians. To explicitly model dynamic objects, we propose an intuitive and effective 4DGS representation of dynamic objects by incorporating a learnable time embedding with each Gaussian and predicting Gaussian deformations at desired timestamps through a multilayer perception (MLP). Moreover, we employ a k-nearest neighbor (KNN)-based consistency regularization mechanism to enforce consistency across neighboring Gaussians to alleviate the inaccurate static reconstruction caused by the low-texture characteristic of the ground surface.
Our extensive experiments on multiple real-world datasets demonstrate that Urban4D achieves comparable or better quality than previous state-of-the-art methods. The proposed framework effectively captures dynamic objects while maintaining high visual fidelity for static elements. By addressing the limitations of existing methods and introducing innovative techniques for semantic-guided decomposition and temporal modeling, Urban4D appears to be a compelling solution for representing the complexity and detail of dynamic urban environments. Our contributions can be summarized as follows:
  1. $\bullet$
    We propose Urban4D, a novel framework that extends the capabilities of 3DGS to effectively handle dynamic urban scenes by leveraging semantic and spatiotemporal information.
  2. $\bullet$
    An intuitive semantic-guided decomposition strategy based on 2D semantic maps, which distinguishes between dynamic and static elements through reliable semantic Gaussians.
  3. $\bullet$
    An intuitive and efficient 4D Gaussian splatting representation that incorporates a learnable time embedding for each Gaussian and predicts deformations over time with an MLP, effectively modeling dynamic objects in urban scenes.
  4. $\bullet$
    A KNN-based consistency regularization mechanism to handle the low-texture ground surface in urban scenes, improving the accuracy of static reconstruction.
2 Related Work
2.1 Neural Scene Representations
Neural representations have revolutionized novel view synthesis, with NeRF [20] leading significant advances in this field. NeRF utilizes multi-layer perceptrons (MLPs) and differentiable volume rendering to reconstruct 3D scenes from 2D images and camera poses. While demonstrating impressive results for bounded scenes, its application to large-scale unbounded scenes remains challenging due to computational constraints and the requirement for consistent camera-object distances.
Various improvements have been proposed to address NeRF’s limitations. Training speed has been enhanced through techniques like voxel grids [10, 25, 9], hash encoding [21], and tensor factorization [6], while rendering quality has been improved through better anti-aliasing [3, 14, 17] and reflection modeling [31, 11].
More recently, 3D Gaussian Splatting [13] has emerged as a promising alternative, offering faster training and rendering while maintaining high quality results. This explicit representation combines the advantages of volumetric rendering with efficient rasterization-based techniques. Compared to previous explicit representations (e.g., mesh, voxels), 3D-GS can model complex shapes while allowing fast, differentiable rendering through splat-based rasterization.
2.2 Dynamic Scene Modeling
Dynamic scene reconstruction methods generally fall into two categories: deformation-based and modulation-based approaches. Deformation-based methods [22, 23, 5, 28, 24] model scene dynamics through canonical space mapping and deformation networks, while modulation-based approaches [16, 15, 33, 18] incorporate temporal information directly. These methods have shown promising results in controlled environments but face significant challenges when applied to complex real-world scenarios with multiple dynamic objects.
For urban environments, several pioneering works have tackled static scene reconstruction by introducing multi-scale NeRF variants [19, 26, 29] and incorporating advanced rendering techniques like mipmapping [1, 2]. Building upon these foundations, recent methods [36, 30, 38] have explored the integration of multi-modal data, combining RGB images with LiDAR point clouds to enhance geometric accuracy. However, the challenge of jointly modeling static and dynamic elements remains complex, particularly due to high-speed movements and sparse viewpoints typical in driving scenarios.
To address these challenges, recent works have proposed various scene decomposition strategies. Scene graph representations [39, 32, 37, 41, 27] enable explicit object-level modeling and control. However, most current approaches either treat all dynamic elements uniformly [7] or rely heavily on manual annotations [41, 37].
3 Methodology
Method overview. Our proposed method aims to reconstruct dynamic urban scenes by leveraging semantic information to effectively distinguish between static and dynamic elements.
Figure 3: Semantic-guided decomposition over time. For each timestamp ($T_{1}$, $T_{2}$, $T_{3}$), semantic Gaussians of the current frame are obtained through rendering and supervision of corresponding semantic maps. Dynamic classes include vehicles, pedestrians, and cyclists, while the static set comprises buildings, vegetation, and roads. For simplicity, we use the ”Road” to represent ground surfaces.
Given a sequence of images $\{I_{t}\}_{t=1}^{T}$ and corresponding LiDAR point clouds $\{P_{t}\}_{t=1}^{T}$ captured by a moving vehicle, we aim to reconstruct dynamic urban scenes. For each frame, semantic maps $\{S_{t}\}_{t=1}^{T}$ are predicted using an off-the-shelf pre-trained segmentation model. Based on our aforementioned insight, we develop a novel framework to intuitively represent dynamic elements in 4D and enhance the reconstruction of static elements.
Our method consists of three main elements (Figure 4): (1) semantic-guided decomposition that separates the scene into static and potentially dynamic Gaussians based on semantic information (Sec. 3.2), (2) 4D Gaussians Splatting representation for dynamic scenes(Sec. 3.3), and (3) a regularization method that enforces consistency in low-texture regions, particularly ground surfaces, through KNN-based optimization (Sec. 3.4).
3.1 Preliminaries
Deformable Gaussian Splatting represents a dynamic scene using a set of deformable 3D Gaussians $\{\mathcal{G}_{i}\}_{i=1}^{N}$. Each Gaussian $\mathcal{G}_{i}$ is characterized by its canonical position $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$, rotation $\mathbf{r}_{i}\in\mathbb{R}^{4}$, and scale $\mathbf{s}_{i}\in\mathbb{R}^{3}$. At time $t$, both position and shape undergo deformation:
$\boldsymbol{\mu}_{i}(t)$ $=\boldsymbol{\mu}_{i}+\mathcal{F}_{pos}(\boldsymbol{\mu}_{i},t),$ (1)
$\mathbf{r}_{i}(t)$ $=\mathbf{r}_{i}+\mathcal{F}_{rot}(\boldsymbol{\mu}_{i},t),$
$\mathbf{s}_{i}(t)$ $=\mathbf{s}_{i}+\mathcal{F}_{scale}(\boldsymbol{\mu}_{i},t),$
where $\mathcal{F}_{pos}$, $\mathcal{F}_{rot}$, and $\mathcal{F}_{scale}$ are deformation MLPs predicting position offset, rotation change, and scale adjustment respectively based on the Gaussian’s position and time. The covariance matrix at time $t$ is then computed as:
$\boldsymbol{\Sigma}_{i}(t)=\mathbf{r}_{i}(t)\mathbf{s}_{i}(t)\mathbf{s}_{i}(t)^{T}\mathbf{r}_{i}(t)^{T}.$ (2)
Each Gaussian also maintains its opacity $\alpha_{i}\in[0,1]$ and color represented as spherical harmonics coefficients $\mathbf{c}_{i}$ for appearance modeling. The rendering process follows standard Gaussian splatting with the deformed parameters.
3.2 Semantic-guided Decomposition
Effectively separating dynamic and static elements is essential for handling urban scenes. Leveraging 2D semantic maps from a pre-trained segmentation model, we introduce a systematic approach to identify and separate potentially dynamic elements, allowing for specialized treatment of movable objects and static elements. Specifically, as shown in Figure 3, for every timestamp, each Gaussian $\mathcal{G}_{i}$ of the current frame is assigned a semantic label $s_{i}\in\{1,...,K\}$, obtained through rendering and supervision of semantic maps, where $K$ represents the total number of semantic classes.
Our method focuses on potentially dynamic objects in urban scenes, typically comprising well-defined semantic categories. We identify dynamic classes $\mathcal{C}_{d}$, including vehicles, pedestrians, and cyclists, which frequently exhibit motion and require specialized handling. The remaining classes form the static set $\mathcal{C}_{s}$, including buildings, vegetation, and road surfaces. This semantic understanding enables the decomposition of the scene into two disjoint sets:
$\mathcal{G}_{i}^{d}=\{\mathcal{G}_{i}|s_{i}\in\mathcal{C}_{d}\},$ (3)
$\mathcal{G}_{i}^{s}=\{\mathcal{G}_{i}|s_{i}\in\mathcal{C}_{s}\},$
where $\mathcal{G}_{i}^{d}$ represents dynamic Gaussians (i.e., Gaussians of potentially dynamic objects) requiring 4D modeling, and $\mathcal{G}_{i}^{s}$ represents static Gaussians. This semantic-guided decomposition offers several key advantages: (1) it enables specialized motion modeling for dynamic objects while preserving static scene consistency, (2) it eliminates the need for labor-intensive manual annotations, and (3) it ensures consistent motion treatment for all dynamic objects, regardless of their movement state.
The decomposition lays the foundation for our two-stream optimization strategy. Dynamic Gaussians in $\mathcal{G}_{i}^{d}$ undergo dedicated motion refinement (Sec. 3.3) to accurately capture movement, while static Gaussians receive geometric regularization (Sec. 3.4) to enhance scene stability. Additionally, we employ an optimizable environment texture map for sky representation, which is rendered separately and composited with the Gaussian-rendered image through alpha blending, as described in [7]. This separation allows targeted optimization without interference between dynamic and static components. Optimizing stationary dynamic Gaussians (e.g., parked vehicles) within a distinct subset is still more efficient than methods like DeformGS [40], which jointly model entire scene dynamics. This targeted approach results in a more accurate reconstruction of both dynamic and static elements.
Figure 4: Overview of Urban4D framework. Given input images with semantic information, Gaussians are classified into static and potentially dynamic elements through semantic-guided decomposition. Dynamic objects are represented in 4D Gaussian Splatting that captures motion patterns by incorporating a learnable time embedding, with deformations predicted at desired timestamps using an MLP. For static Gaussians, particularly low-texture areas like ground surfaces, consistency is enforced using a KNN-based regularization mechanism.
3.3 4D Gaussian Splatting Representation
To represent our dynamic Gaussians $\{\mathcal{G}_{i}^{d}\}$ in 4D, our method extends the deformation mechanism of DeformGS [40] by introducing a learnable time embedding for each Gaussian. While DeformGS [40] directly maps spatial positions and time to deformations, our approach leverages temporal context through Gaussian-specific embeddings.
Specifically, for each dynamic Gaussian $\mathcal{G}_{i}^{d}$, we maintain a learnable time embedding vector $\mathbf{e}_{i}\in\mathbb{R}^{D_{e}}$. At time step $t$, we form the input feature by concatenating this temporal embedding with position and time information:
$\mathbf{h}_{i}(t)=[\boldsymbol{\mu}_{i};t;\mathbf{e}_{i}],$ (4)
where [;][;][ ; ] denotes concatenation, $t$ is the normalized time stamp, and $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$ represents the original 3D position of the Gaussian. This temporal-aware design enables more accurate modeling of complex motions compared to the direct mapping used in DeformGS [40].
This combined feature vector is processed by a lightweight MLP to predict residual corrections:
$[\Delta\boldsymbol{\mu}_{i}(t),\Delta\alpha_{i}(t),\Delta\boldsymbol{r}_{i}(t),\Delta\boldsymbol{s}_{i}(t)]=\text{MLP}(\mathbf{h}_{i}(t)),$ (5)
where $\Delta\boldsymbol{\mu}_{i}(t)\in\mathbb{R}^{3}$, $\Delta\alpha_{i}(t)\in\mathbb{R}$, $\Delta\boldsymbol{r}_{i}(t)\in\mathbb{R}^{4}$, $\Delta\boldsymbol{s}_{i}(t)\in\mathbb{R}^{3}$ are the predicted position, opacity, rotation and scale residuals, respectively. The final parameters of 4D Gaussians at time $t$ are obtained by:
$\boldsymbol{\mu}_{i}^{\prime}(t)$ $=\boldsymbol{\mu}_{i}+\Delta\boldsymbol{\mu}_{i}(t),$ (6)
$\alpha_{i}^{\prime}(t)$ $=\alpha_{i}+\Delta\alpha_{i}(t),$
$\boldsymbol{r}_{i}^{\prime}(t)$ $=\boldsymbol{r}_{i}+\Delta\boldsymbol{r}_{i}(t),$
$\boldsymbol{s}_{i}^{\prime}(t)$ $=\boldsymbol{s}_{i}+\Delta\boldsymbol{s}_{i}(t),$
where $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$, $\alpha_{i}\in[0,1]$, $\boldsymbol{r}_{i}\in\mathbb{R}^{3}$,$\boldsymbol{s}_{i}\in\mathbb{R}^{3}$ denotes the initial 3D position, opacity, rotation and scaling of the $i$-th Gaussian. These refined parameters are then used in the standard 3D Gaussian Splatting rendering process to generate the final images. This refinement mechanism allows each dynamic Gaussians to adapt its attributions based on its temporal context, enabling more accurate representation of moving objects in the scene.
The MLP architecture is intentionally kept lightweight to maintain computational efficiency while providing sufficient capacity for modeling temporal dynamics. The detailed architecture of MLPs used in our method is provided in the supplementary material for reproducibility. The entire refinement process is end-to-end trainable along with the main 3DGS optimization objectives.
3.4 Ground Surface Consistency Regularization
In urban driving scenes, ground surfaces constitute a significant portion of the environment and typically exhibit low-texture characteristics. While ground-level Gaussians should theoretically share similar properties due to their homogeneous nature, enforcing strict uniformity across all ground Gaussians would be oversimplified and impractical, as real-world surfaces often contain variations and irregularities. The scale parameter of a Gaussian, derived from its covariance matrix, inherently encodes local geometric information analogous to surface normals [12, 8]. A well-behaved ground surface should exhibit smooth transitions in its local geometry, making scale a particularly suitable target for regularization. This motivates us to regularize the scale parameters rather than other Gaussian properties.
For each ground Gaussian $\mathcal{G}_{i}\in\mathcal{G}_{g}$ (where $\mathcal{G}_{g}\subset\mathcal{G}_{i}^{s}$ denotes the set of ground surface Gaussians), we identify its $N$ nearest neighbors:
$\mathcal{N}_{i}=\text{KNN}(\mathcal{G}_{i},\mathcal{G}_{g},N),$ (7)
where KNN retrieves the $N$ spatially closest Gaussians to $\mathcal{G}_{i}$ from the global set $\mathcal{G}_{g}$. The neighbors are determined based on the Euclidean distance between Gaussian centers $\boldsymbol{\mu}_{i}$, forming a local neighborhood for geometric consistency. We then introduce a local consistency loss that encourages similar scale properties within each local neighborhood:
$\mathcal{L}_{\text{ground}}=\sum_{\mathcal{G}_{i}\in\mathcal{G}_{g}}\|\boldsymbol{s}_{i}-\frac{1}{N}\sum_{\mathcal{G}_{j}\in\mathcal{N}_{i}}\boldsymbol{s}_{j}\|_{2}^{2},$ (8)
where $\boldsymbol{s}_{i}$ and $\boldsymbol{s}_{j}$ represents the scale parameter of the Gaussian $\mathcal{G}_{i}$ and $\mathcal{G}_{j}$. By regularizing the scale parameters, we effectively enforce consistency in the local surface geometry while preserving the ability to model natural surface variations. This approach leads to more coherent ground surface reconstruction, as similar scale parameters in a local neighborhood implicitly enforce consistent surface normal orientations, resulting in improved geometric fidelity of the ground surface representation.
3.5 Optimization Strategy
Our optimization objective comprises multiple loss terms that jointly ensure high-quality visual rendering, geometric accuracy, and semantic consistency. The overall loss function is formulated as follows:
$\mathcal{L}_{\text{total}}=$ $\lambda_{1}\mathcal{L}_{\text{L1}}+\lambda_{2}\mathcal{L}_{\text{SSIM}}+\lambda_{3}\mathcal{L}_{\text{sem}}$
$+\lambda_{4}\mathcal{L}_{\text{ground}}+\lambda_{5}\mathcal{L}_{\text{depth}}+\lambda_{6}\mathcal{L}_{\text{sky}}$ (9)
where $\{\lambda_{i}\}_{i=1}^{6}$ are weighting coefficients balancing different loss terms. Each loss term serves a specific purpose in our optimization.
Appearance Supervision. The L1 loss and SSIM loss work together to ensure accurate color reproduction and structural similarity:
$\mathcal{L}_{\text{L1}}$ $=\|I_{\text{rendered}}-I_{\text{gt}}\|_{1},$ (10)
$\mathcal{L}_{\text{SSIM}}$ $=1-\text{SSIM}(I_{\text{rendered}},I_{\text{gt}}),$
where $I_{\text{rendered}}$ and $I_{\text{gt}}$ denote the rendered image and ground-truth image respectively.
Semantic Consistency. The semantic loss ensures correct class predictions for each Gaussian:
$\mathcal{L}_{\text{sem}}=\text{CE}(S_{\text{rendered}},S_{\text{gt}}),$ (11)
where CE denotes cross-entropy loss between rendered semantic maps $S_{\text{rendered}}$ and ground-truth semantic maps $S_{\text{gt}}$.
Ground Consistency. As detailed in Sec. 3.4, the ground consistency loss enforces local geometric similarity as defined in Eq. 8.
Geometric Supervision. The inverse depth loss aligns the scene geometry with LiDAR measurements:
$\mathcal{L}_{\text{depth}}=\|\frac{1}{D_{\text{rendered}}}-\frac{1}{D_{\text{lidar}}}\|_{1},$ (12)
where $D_{\text{rendered}}$ and $D_{\text{lidar}}$ represent the rendered depth and LiDAR depth respectively.
Sky Region Handling. For sky regions, we encourage low opacity to prevent incorrect geometry:
$\mathcal{L}_{\text{sky}}=\sum_{\mathcal{G}_{i}\in\mathcal{G}_{\text{sky}}}\|\alpha_{i}\|_{1},$ (13)
where $\mathcal{G}_{\text{sky}}$ represents the set of sky Gaussians.
These joint loss terms collectively constrain the scene reconstruction process, ensuring high-quality results.
Figure 5: comparison of reconstruction quality across consecutive frames. DeformGS [40] struggles significantly with reconstructing dynamic objects, resulting in severe artifacts and a failure to accurately represent motion. PVG [7] captures dynamic vehicles to some extent but suffers from noticeable blurring, particularly in the lower parts of the objects. In contrast, Urban4D delivers superior reconstruction quality, maintaining high fidelity and preserving clear details throughout the dynamic objects.
Table 1: Quantitative comparison with existing methods on the nuScenes [4] dataset. The best results in each group are highlighted in bold. Our method outperforms previous approaches across all evaluation metrics.
Method Full Image Non-Sky Human Vehicle Dynamic
[…]
SUDS [30] 20.02 0.605 20.45 0.621 21.20 0.633 21.98 0.678 21.49 0.646
DeformGS [40] 25.32 0.825 25.27 0.822 21.44 0.6456 21.72 0.700 21.62 0.684
3DGS [13] 26.02 0.825 26.45 0.836 23.20 0.721 23.98 0.794 23.59 0.756
EmerNeRF [38] 26.12 0.830 26.50 0.840 23.45 0.733 24.69 0.808 24.12 0.767
PVG [7] 26.23 0.834 26.72 0.841 23.98 0.743 24.73 0.815 24.33 0.774
Urban4D (Ours) 26.47 0.839 28.93 0.878 24.88 0.783 26.05 0.855 25.33 0.810
4 Experiments
4.1 Implementation Details
We initialize Gaussian points from both LiDAR points (with projected RGB and semantic values) and 200K random points sampled within a sphere. Following previous 3DGS-based methods to predict reliable semantic Gaussians, we use SegFormer [35] as our pre-trained segmentation model. The learning rate of MLP starts from $1.6\times 10^{-4}$ and decreases to $1.6\times 10^{-6}$. For each loss term, the weighting coefficients are empirically set to $\lambda_{1}=0.8$, $\lambda_{2}=0.2$, $\lambda_{3}=0.01$, $\lambda_{4}=0.0001$, $\lambda_{5}=0.1$ and $\lambda_{6}=0.01$. All experiments are conducted on a single NVIDIA V100s with 32GB memory.
4.2 Datasets
Our experiments are conducted on two popular autonomous driving datasets: PandaSet [7] and nuScenes [4]. We primarily evaluate our method on nuScenes [4], which consists of 1000 scenes captured in Boston and Singapore under diverse urban scenarios with multimodal sensor data, featuring challenging weather conditions and complex urban scenes. We perform ablation studies on PandaSet [7], which is a comprehensive autonomous driving dataset captured in San Francisco, containing 103 sequences of synchronized LiDAR-camera data and images. For both datasets, we use the officially provided camera poses.
4.3 Results and Comparisons
Results on nuScenes [4]. We evaluate our method against previous approaches on the nuScenes [4] dataset, with a comprehensive comparison across different scene components and object categories. As shown in Table 1, our method demonstrates consistent improvements over NeRF-based methods [30, 38], traditional 3DGS [13], PVG [7] and DeformGS [40] across all evaluation metrics, with significant improvements in dynamic object reconstruction.
For full image reconstruction, our method achieves 26.47 PSNR and 0.839 SSIM, surpassing PVG [7] by 0.24 PSNR. The improvement is more pronounced in non-sky regions, where our approach shows a substantial gain of 2.21 PSNR (28.93 vs 26.72) and 0.037 SSIM (0.878 vs 0.841) compared to PVG [7], indicating better handling of complex geometric structures in urban environments.
The advantages of our method are most evident in dynamic object reconstruction. For human instances, we achieve 24.88 PSNR and 0.783 SSIM, outperforming PVG [7] by 0.90 PSNR and 0.040 SSIM. In vehicle reconstruction, our method demonstrates even larger improvements, reaching 26.05 PSNR and 0.855 SSIM, surpassing PVG [7] by 1.32 PSNR and 0.040 SSIM. These significant gains can be attributed to our dynamic-focused semantic decomposition strategy, which enables specialized treatment of moving objects.
We present qualitative comparisons in Fig. 5. For dynamic objects like vehicles and pedestrians, our method shows significant improvements over PVG [7] and DeformGS [40], with notably reduced motion blur and sharper object boundaries. This is particularly evident in regions with fast-moving objects, where previous methods tend to produce blurry reconstructions. Meanwhile, our approach also demonstrates better reconstruction quality for static scene elements such as roads, preserving more detailed textures and geometric structures. These visual improvements align well with our quantitative results, where we achieve consistently higher scores across both dynamic and static regions. The qualitative results in Fig. 5 further support these findings. In consecutive frame reconstruction, DeformGS [40] fails to handle dynamic objects, producing severe artifacts. While PVG [7] captures the overall shape of moving vehicles, it suffers from noticeable blurring artifacts, particularly in the lower parts of the vehicles. In contrast, our method achieves clearer and more consistent reconstruction across all dynamic objects.
Table 2: Quantitative comparison with state-of-the-art methods on the PandaSet [34] dataset. We report image reconstruction and novel view synthesis metrics. The best results are shown in bold.
Method Image Reconstruction Novel View Synthesis
[…]
3DGS [13] 23.67 0.743 22.14 0.713
EmerNeRF [38] 26.45 0.812 24.89 0.765
PVG [7] 27.15 0.836 25.92 0.798
Urban4D (Ours) 27.74 0.850 26.56 0.814
Results on PandaSet [34]. To further validate the generalization capability of our method, we conduct additional experiments on the PandaSet [34] dataset. As shown in Table 2, our approach demonstrates consistently superior performance across both image reconstruction and novel view synthesis tasks. For image reconstruction, our method achieves 27.74 PSNR and 0.850 SSIM, outperforming PVG [7] by 0.59 PSNR and 0.014 SSIM. The performance advantage is even more pronounced when compared to vanilla 3D Gaussian Splatting (3DGS [13]), showing improvements of 4.07 PSNR and 0.107 SSIM. This significant gap highlights the effectiveness of our dynamic-focused scene decomposition strategy.
In the more challenging novel view synthesis task, our method maintains its lead with 26.56 PSNR and 0.814 SSIM, surpassing PVG [7] by 0.64 PSNR and 0.016 SSIM. Notably, while all methods experience some performance degradation in novel view synthesis compared to image reconstruction, our method shows the smallest drop (1.18 PSNR vs 1.23 PSNR for PVG [7]), indicating better scene understanding and geometry modeling. The consistent performance improvements across different datasets and evaluation metrics demonstrate the robustness and generalizability of our approach across varied urban scenes and capture conditions.
4.4 Ablation Study
Effectiveness of consistency regularization. To validate the effectiveness of our ground surface consistency module, we conduct extensive experiments on the PandaSet [34] dataset. Table 3 shows the quantitative results of different variants of our method. The baseline model without road consistency achieves 26.31 PSNR and 0.810 SSIM. After incorporating our consistency constraint, the performance significantly improves by 0.25 dB in PSNR (26.56 vs. 26.31) and 0.004 in SSIM (0.814 vs. 0.810). This demonstrates that enforcing local geometric and appearance consistency on ground surfaces effectively enhances the reconstruction quality.
We further investigate the impact of different numbers of nearest neighbors in our module. As shown in Table 3, using 16 neighbors yields the best performance (26.56 PSNR), outperforming both smaller (k=8, 26.45 PSNR) and larger (k=32, 26.48 PSNR) neighbor sizes. We hypothesize that 8 neighbors are insufficient to capture the local surface structure, while 32 neighbors may introduce noise from points that are too far away. The optimal choice of 16 neighbors provides a good local structure preservation.
Table 3: Ablation study on road consistency module. We evaluate different design choices and parameter settings of our road consistency constraint on PandaSet [7].
[…]
Baseline (w/o road consistency) 26.31 0.810
KNN-8 26.45 0.812
KNN-16 (Ours) 26.56 0.814
KNN-32 26.48 0.812
Table 4: Ablation study on time embedding. We report PSNR and SSIM on the PandaSet [34].
[…]
w/o Time Embeddings 26.40 0.810
w/ Time Embeddings 26.56 0.814
Effectiveness of Time Embeddings. We also validate the effectiveness of our temporal modeling strategy through ablation studies. As shown in Table 4, equipping each Gaussian with a learnable time embedding improves the rendering quality (PSNR: 26.40 → 26.56, SSIM: 0.810 → 0.814) on the PandaSet [34] dataset. This demonstrates that point-specific temporal context helps capture more accurate deformation patterns compared to directly mapping spatial positions to deformations.
5 Conclusions
In conclusion, Urban4D provides a novel semantic-guided decomposition and 4D representation strategy for reconstructing dynamic urban scenes. By leveraging 2D semantic information, our approach effectively separates static and potentially dynamic components. We introduced specialized processing methods for different elements: a 4D Gaussian Splatting representation for dynamic objects and KNN-based consistency regularization for the ground surface. Both quality and quantitative results demonstrated that Urban4D improves rendering quality across various scene components, particularly for dynamic objects and low-texture regions.

### v2

@@
-Urban4D: Semantic-Guided 4D Gaussian Splatting for Urban Scene Reconstruction
+UrbanGS: Semantic-Guided Gaussian Splatting for Urban Scene Reconstruction
 Abstract
-Reconstructing dynamic urban scenes presents significant challenges due to their intrinsic geometric structures and spatiotemporal dynamics.
-Existing methods that attempt to model dynamic urban scenes without leveraging priors on potentially moving regions often produce suboptimal results.
-Meanwhile, approaches based on manual 3D annotations yield improved reconstruction quality but are impractical due to labor-intensive labeling.
-In this paper, we revisit the potential of 2D semantic maps for classifying dynamic and static Gaussians and integrating spatial and temporal dimensions for urban scene representation.
-We introduce Urban4D, a novel framework that employs a semantic-guided decomposition strategy inspired by advances in deep 2D semantic map generation.
-Our approach distinguishes potentially dynamic objects through reliable semantic Gaussians.
-To explicitly model dynamic objects, we propose an intuitive and effective 4D Gaussian splatting (4DGS) representation that aggregates temporal information through learnable time embeddings for each Gaussian, predicting their deformations at desired timestamps using a multilayer perceptron (MLP).
-For more accurate static reconstruction, we also design a k-nearest neighbor (KNN)-based consistency regularization to handle the ground surface due to its low-texture characteristic.
-Extensive experiments on real-world datasets demonstrate that Urban4D not only achieves comparable or better quality than previous state-of-the-art methods but also effectively captures dynamic objects while maintaining high visual fidelity for static elements.
+Reconstructing urban scenes is challenging due to their complex geometries and the presence of potentially dynamic objects.
+3D Gaussian Splatting (3DGS)-based methods have shown strong performance, but existing approaches often incorporate manual 3D annotations to improve dynamic object modeling, which is impractical due to high labeling costs.
+Some methods leverage 4D Gaussian Splatting (4DGS) to represent the entire scene, but they treat static and dynamic objects uniformly, leading to unnecessary updates for static elements and ultimately degrading reconstruction quality.
+To address these issues, we propose UrbanGS, which leverages 2D semantic maps and an existing dynamic Gaussian approach to distinguish static objects from the scene, enabling separate processing of definite static and potentially dynamic elements.
+Specifically, for definite static regions, we enforce global consistency to prevent unintended changes in dynamic Gaussian and introduce a K-nearest neighbor (KNN)-based regularization to improve local coherence on low-textured ground surfaces.
+Notably, for potentially dynamic objects, we aggregate temporal information using learnable time embeddings, allowing each Gaussian to model deformations over time.
+Extensive experiments on real-world datasets demonstrate that our approach outperforms state-of-the-art methods in reconstruction quality and efficiency, accurately preserving static content while capturing dynamic elements.
 Figure 1: Qualitative comparison on the nuScenes [4] dataset.
@@
 In contrast, our Urban4D maintains high fidelity for both dynamic objects and static backgrounds, also surpassing the reconstruction quality of PVG [7].
-Figure 2: Different from previous 3DGS-based methods which leverage pre-trained segmentation models to predict 2D semantic maps for downstream tasks, we revisit the potential of these estimated semantic maps to handle potentially dynamic objects for reconstruction.
-From left to right: input image, semantic map, our reconstruction, and reconstruction without potentially dynamic objects.
-Our method effectively preserves the geometry and appearance of both static structures and potentially dynamic objects.
 1 Introduction
-Recent advancements in urban scene reconstruction have witnessed remarkable progress, particularly with the introduction of 3D Gaussian Splatting (3DGS)  [13], which has enhanced real-time rendering while maintaining high reconstruction quality for static scenes.
-However, directly extending these capabilities to model dynamic and large-scale urban scenes presents significant challenges due to spatiotemporal dynamics and intricate geometric structures.
-Urban scenes can generally be categorized into two types: dynamic objects, such as vehicles and pedestrians, which exhibit complex motion patterns, and static elements, which maintain consistent spatial relationships.
-Specifically, low-texture regions (e.g., ground surfaces) pose challenges for accurate static reconstruction due to their lack of salient patterns.
-These unique characteristics make reconstructing dynamic objects in urban scenes particularly difficult.
-The coexistence of static and dynamic elements necessitates a model that can distinguish between and process them independently to ensure accurate reconstruction.
-Previous dynamic NeRF-based methods, such as  [38, 32], attempt to address these challenges by introducing time-conditioned neural networks to model motion, but they suffer from slow rendering speeds and limited capacity to handle complex and large-scale urban scenes.
-Some methods explore suitable representations, such as Periodic Vibration Gaussian (PVG)  [7], which proposes a unified framework by incorporating periodic vibration-based temporal dynamics in urban scenes.
-However, such a unified representation may degrade performance when treating static elements with temporal processes.
-Other 3DGS-based methods resort to manually labeled 3D bounding boxes to process dynamic objects  [37, 41], which is labor-intensive and impractical for large-scale or continuously changing environments.
-Furthermore, we have witnessed that some of these methods adopt pre-trained segmentation models, such as SegFormer [35], for semantic novel view rendering.
-Semantic segmentation based on deep learning provides more reliable classifications of scene components in 2D images.
-Leveraging these results, we may effectively classify dynamic and static Gaussians without relying on manually labeled 3D annotations.
-The observations in Figure 2 lead us to an intriguing question: Can we revisit the potential of 2D semantic maps for classifying potentially dynamic objects and leverage spatiotemporal information to adequately represent urban scenes?
-The answer to this question uncovers an unexplored area in reconstructing dynamic objects based on 4D Gaussian splatting (4DGS), enabling a more holistic and precise representation.
-Inspired by this perspective, we propose Urban4D, a novel framework to overcome the limitations of traditional 3DGS-based methods in handling temporal dynamics and reconstructing urban scenes.
-Specifically, we design a semantic-guided decomposition strategy, which builds on recent progress in deep 2D semantic map generation to distinguish potentially dynamic objects from static elements by assigning reliable semantic classes to Gaussians.
-To explicitly model dynamic objects, we propose an intuitive and effective 4DGS representation of dynamic objects by incorporating a learnable time embedding with each Gaussian and predicting Gaussian deformations at desired timestamps through a multilayer perception (MLP).
-Moreover, we employ a k-nearest neighbor (KNN)-based consistency regularization mechanism to enforce consistency across neighboring Gaussians to alleviate the inaccurate static reconstruction caused by the low-texture characteristic of the ground surface.
-Our extensive experiments on multiple real-world datasets demonstrate that Urban4D achieves comparable or better quality than previous state-of-the-art methods.
-The proposed framework effectively captures dynamic objects while maintaining high visual fidelity for static elements.
-By addressing the limitations of existing methods and introducing innovative techniques for semantic-guided decomposition and temporal modeling, Urban4D appears to be a compelling solution for representing the complexity and detail of dynamic urban environments.
-Our contributions can be summarized as follows:
+Urban scenes are characterized by two primary categories of objects: major static elements, including buildings and road infrastructure, which remain spatially consistent over time, and some potentially dynamic elements, such as pedestrians and vehicles, which can remain static or exhibit diverse and often unpredictable motion patterns.
+Accurate reconstruction of urban scenes thus remains challenging, mainly due to the coexistence of these static and dynamic elements and complexities arising from low-textured regions.
+Recent advancements in 3D Gaussian Splatting (3DGS)  [43, 48, 8, 10], which have attempted to incorporate manually labeled 3D bounding boxes to process dynamic objects separately.
+However, such manual annotations are labor-intensive, impractical for large-scale settings, and unsuitable for continuously evolving environments where dynamic objects frequently change positions.
+Alternative approaches leverage 4DGS-based representations, such as Periodic Vibration Gaussian (PVG) [7], which introduces periodic temporal modeling to represent motion variations in urban scenes.
+However, these methods lack explicit differentiation between static and dynamic elements, leading to unnecessary updates for stationary objects, which in turn degrades reconstruction quality.
+We observe that Gaussians supervised by semantic maps inherently acquire semantic information, which can be leveraged for identifying static objects.
+Furthermore, deep-learning-based 2D semantic segmentation models provide robust classification capabilities, allowing Gaussians to be categorized into determined static and potentially dynamic elements without relying on explicit 3D annotations.
+Inspired by this perspective, we introduce UrbanGS, a semantic-guided Gaussian Splatting framework designed to effectively handle definite static elements in urban scene reconstruction while adapting to potentially dynamic components.
+Specifically, for definite static Gaussians, we introduce a global consistency constraint to ensure that they remain unchanged over time.
+Additionally, we employ a K-nearest neighbor (KNN)-based consistency regularization to improve local coherence, particularly for low-textured surfaces such as roads, which pose significant challenges in urban scene reconstruction.
+For potentially dynamic Gaussians, we propose an efficient 4DGS representation that incorporates learnable time embeddings for each Gaussian.
+This design enables the model to predict object deformations at arbitrary timestamps using a lightweight multilayer perceptron (MLP), effectively capturing urban dynamics while preserving rendering efficiency.
+Our extensive experiments on real-world urban datasets demonstrate that UrbanGS achieves state-of-the-art reconstruction quality in both static and potentially dynamic objects.
+The key contributions of this work can be summarized as follows:
   1. $\bullet$
-    We propose Urban4D, a novel framework that extends the capabilities of 3DGS to effectively handle dynamic urban scenes by leveraging semantic and spatiotemporal information.
+    We introduce UrbanGS, a novel semantic-driven framework that leverage 2D semantic segmentation to separate static Gaussians from potentially dynamic Gaussians without requiring manual 3D annotations.
   2. $\bullet$
-    An intuitive semantic-guided decomposition strategy based on 2D semantic maps, which distinguishes between dynamic and static elements through reliable semantic Gaussians.
+    We propose a global consistency constraint to enforce temporal stability in static Gaussians, preventing unnecessary updates and significantly improving reconstruction quality.
+Additionally, to address the challenge of low-textured regions, we introduce a KNN-based consistency regularization, ensuring a more stable and accurate reconstruction of surfaces such as roads and sidewalks.
   3. $\bullet$
-    An intuitive and efficient 4D Gaussian splatting representation that incorporates a learnable time embedding for each Gaussian and predicts deformations over time with an MLP, effectively modeling dynamic objects in urban scenes.
-  4. $\bullet$
-    A KNN-based consistency regularization mechanism to handle the low-texture ground surface in urban scenes, improving the accuracy of static reconstruction.
+    We develop a learnable time embedding mechanism for potentially dynamic Gaussians, enabling the model to predict object deformations at arbitrary timestamps using a lightweight MLP-based deformation model, efficiently handling motion in urban scenes.
+Figure 2: Semantic-guided decomposition over time.
+For each timestamp ($T_{1}$, $T_{2}$, $T_{3}$), semantic Gaussians of the current frame are obtained through rendering and supervision of corresponding semantic maps.
+Dynamic classes include vehicles, pedestrians, and cyclists, while the static set comprises buildings, vegetation, and roads.
+For simplicity, we use the ”Road” to represent ground surfaces.
 2 Related Work
-2.1 Neural Scene Representations
-Neural representations have revolutionized novel view synthesis, with NeRF [20] leading significant advances in this field.
+Neural Scene representations have revolutionized novel view synthesis, with NeRF [24] leading significant advances in this field.
 NeRF utilizes multi-layer perceptrons (MLPs) and differentiable volume rendering to reconstruct 3D scenes from 2D images and camera poses.
@@
 Compared to previous explicit representations (e.g., mesh, voxels), 3D-GS can model complex shapes while allowing fast, differentiable rendering through splat-based rasterization.
-2.2 Dynamic Scene Modeling
 Dynamic scene reconstruction methods generally fall into two categories: deformation-based and modulation-based approaches.
@@
 To address these challenges, recent works have proposed various scene decomposition strategies.
-Scene graph representations [39, 32, 37, 41, 27] enable explicit object-level modeling and control.
-However, most current approaches either treat all dynamic elements uniformly [7] or rely heavily on manual annotations [41, 37].
+Scene graph representations [45, 38, 43, 48, 32, 8, 10] enable explicit modeling and control at the object level.
+However, most current approaches either treat all dynamic elements uniformly [7, 46, 15] or rely heavily on manual annotations [48, 43, 8, 10].
+Gaussian reconstruction with semantic features Recent advances integrate 3DGS with semantic features.
+Feature 3DGS [47] extends 3D Gaussian Splatting by introducing high-dimensional feature fields.
+Similarly, Semantic Gaussians [13] tackles open-vocabulary 3D scene understanding by mapping diverse 2D semantic features into 3D Gaussian.
+These works focus on static scenes, while our work complements these efforts by focusing on urban scenes, leveraging semantic decomposition for static/dynamic separation.
 3 Methodology
-Method overview. Our proposed method aims to reconstruct dynamic urban scenes by leveraging semantic information to effectively distinguish between static and dynamic elements.
-Figure 3: Semantic-guided decomposition over time.
-For each timestamp ($T_{1}$, $T_{2}$, $T_{3}$), semantic Gaussians of the current frame are obtained through rendering and supervision of corresponding semantic maps.
-Dynamic classes include vehicles, pedestrians, and cyclists, while the static set comprises buildings, vegetation, and roads.
-For simplicity, we use the ”Road” to represent ground surfaces.
-Given a sequence of images $\{I_{t}\}_{t=1}^{T}$ and corresponding LiDAR point clouds $\{P_{t}\}_{t=1}^{T}$ captured by a moving vehicle, we aim to reconstruct dynamic urban scenes.
+Figure 3: Overview of UrbanGS framework.
+Given input images with semantic information during training, Gaussians are classified into definite static and potentially dynamic elements through semantic-guided decomposition.
+For definitively static Gaussians, we introduce a static invariance constraint to preserve their temporal invariance and prevent unintended transformations.
+To address challenges in low-texture regions (e.g., ground surfaces), a KNN-based regularization mechanism is employed to enforce structural coherence.
+Potentially dynamic objects are represented in 4D Gaussian Splatting that captures motion patterns by incorporating a learnable time embedding, with deformations predicted at desired timestamps using an MLP.
+Method overview. Our proposed method aims to reconstruct dynamic urban scenes by leveraging semantic information to effectively distinguish between determined static and potentially dynamic elements.
+Given a sequence of images $\{I_{t}\}_{t=1}^{T}$ and the corresponding LiDAR point clouds $\{P_{t}\}_{t=1}^{T}$ captured by a moving vehicle, our aim is to reconstruct urban scenes.
 For each frame, semantic maps $\{S_{t}\}_{t=1}^{T}$ are predicted using an off-the-shelf pre-trained segmentation model.
-Based on our aforementioned insight, we develop a novel framework to intuitively represent dynamic elements in 4D and enhance the reconstruction of static elements.
+Building upon a uniformly dynamic Gaussian approach, which inherently applies time-dependent transformations to the global Gaussians, such as PVG [7], we propose a novel framework that introduces semantic-aware improvement.
+Specifically, during training, we leverage the semantic attributes of each Gaussian to enforce constraints and adaptively adjust the properties of each Gaussian point.
+This allows us to effectively keep static elements unchanged across time, enforce consistency in low-texture regions, and intuitively capture potentially dynamic objects through a 4D representation.
+Our approach enhances the robustness and accuracy of urban scene reconstruction by combining the strengths of semantic guidance and dynamic Gaussian modeling.
+More details of pseudo algorithm for training are included in the supplementary material.
 Our method consists of three main elements (Figure 3): (1) semantic-guided decomposition that separates the scene into static and potentially dynamic Gaussians based on semantic information (Sec.
-3.2), (2) 4D Gaussians Splatting representation for dynamic scenes(Sec.
-3.3), and (3) a regularization method that enforces consistency in low-texture regions, particularly ground surfaces, through KNN-based optimization (Sec.
+3.2), (2) a temporal-invariance regularization for all static points, ensuring they remain unchanged over time, and apply a KNN-based consistency constraint to low-texture regions for enhanced reconstruction fidelity (Sec.
+3.3), and (3) a 4d Gaussians Splatting representation for potentially dynamic objects (Sec.
 3.4) .
 3.1 Preliminaries
-Deformable Gaussian Splatting represents a dynamic scene using a set of deformable 3D Gaussians $\{\mathcal{G}_{i}\}_{i=1}^{N}$.
-Each Gaussian $\mathcal{G}_{i}$ is characterized by its canonical position $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$, rotation $\mathbf{r}_{i}\in\mathbb{R}^{4}$, and scale $\mathbf{s}_{i}\in\mathbb{R}^{3}$.
-At time $t$, both position and shape undergo deformation:
-$\boldsymbol{\mu}_{i}(t)$ $=\boldsymbol{\mu}_{i}+\mathcal{F}_{pos}(\boldsymbol{\mu}_{i},t),$ (1)
-$\mathbf{r}_{i}(t)$ $=\mathbf{r}_{i}+\mathcal{F}_{rot}(\boldsymbol{\mu}_{i},t),$
-$\mathbf{s}_{i}(t)$ $=\mathbf{s}_{i}+\mathcal{F}_{scale}(\boldsymbol{\mu}_{i},t),$
-where $\mathcal{F}_{pos}$, $\mathcal{F}_{rot}$, and $\mathcal{F}_{scale}$ are deformation MLPs predicting position offset, rotation change, and scale adjustment respectively based on the Gaussian’s position and time.
-The covariance matrix at time $t$ is then computed as:
-$\boldsymbol{\Sigma}_{i}(t)=\mathbf{r}_{i}(t)\mathbf{s}_{i}(t)\mathbf{s}_{i}(t)^{T}\mathbf{r}_{i}(t)^{T}.$ (2)
-Each Gaussian also maintains its opacity $\alpha_{i}\in[0,1]$ and color represented as spherical harmonics coefficients $\mathbf{c}_{i}$ for appearance modeling.
-The rendering process follows standard Gaussian splatting with the deformed parameters.
+3D Gaussian Splatting represents a scene as a set of 3D Gaussians $\{\mathcal{G}_{i}\}_{i=1}^{N}$, where each Gaussian $\mathcal{G}_{i}$ is parameterized by its mean position $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$, covariance matrix $\boldsymbol{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$, and appearance features including opacity $\alpha_{i}\in\mathbb{R}$ and spherical harmonics coefficients $\mathbf{f}_{i}\in\mathbb{R}^{48}$ for RGB color representation.
+For each pixel in the target view, the rendering process involves projecting the 3D Gaussians onto the 2D image plane. The projection of a 3D Gaussian results in a 2D Gaussian with parameters:
+$\boldsymbol{\mu}_{2D}=\Pi(\boldsymbol{\mu}_{i}),$ (1)
+$\boldsymbol{\Sigma}_{2D}=J\boldsymbol{\Sigma}_{i}J^{T},$ (2)
+where $\Pi(\cdot)$ is the perspective projection function and $J$ is the Jacobian of the projection.
+The final color $C(x,y)$ at pixel $(x,y)$ is computed through alpha compositing:
+$C(x,y)=\sum_{i=1}^{N}T_{i}\alpha_{i}c_{i},$ (3)
+where $T_{i}$ represents the accumulated transmittance, $\alpha_{i}$ is the opacity, and $c_{i}$ is the RGB color from spherical harmonics coefficients.
+The scene is optimized by minimizing the difference between rendered and ground-truth images.
+In addition to RGB supervision, each Gaussian can be associated with semantic labels $s_{i}\in\{1,...,K\}$, where $K$ is the number of semantic classes.
+The semantic prediction at pixel $(x,y)$ can be computed similarly through alpha compositing:
+$S(x,y)=\sum_{i=1}^{N}T_{i}\alpha_{i}s_{i},$ (4)
 3.2 Semantic-guided Decomposition
-Effectively separating dynamic and static elements is essential for handling urban scenes.
-Leveraging 2D semantic maps from a pre-trained segmentation model, we introduce a systematic approach to identify and separate potentially dynamic elements, allowing for specialized treatment of movable objects and static elements.
-Specifically, as shown in Figure 3, for every timestamp, each Gaussian $\mathcal{G}_{i}$ of the current frame is assigned a semantic label $s_{i}\in\{1,...,K\}$, obtained through rendering and supervision of semantic maps, where $K$ represents the total number of semantic classes.
-Our method focuses on potentially dynamic objects in urban scenes, typically comprising well-defined semantic categories.
-We identify dynamic classes $\mathcal{C}_{d}$, including vehicles, pedestrians, and cyclists, which frequently exhibit motion and require specialized handling.
-The remaining classes form the static set $\mathcal{C}_{s}$, including buildings, vegetation, and road surfaces.
+Leveraging 2D semantic maps from a pre-trained segmentation model, we introduce a systematic approach to identify and separate static elements, allowing for specialized treatment of static elements and remaining potentially movable objects.
+Specifically, as shown in Figure 2, during dynamic Gaussian training, each Gaussian point $\mathcal{G}_{i}$ is assigned a semantic label $s_{i}\in\{1,...,K\}$, obtained through rendering and supervision of semantic maps, where $K$ represents the total number of semantic classes.
+Our method focuses on determined static objects in urban scenes, typically comprising well-defined semantic categories such as building, roads and trees.
+We have designed a strategy specifically for those static objects to ensure their invariance over time.
+We also identify potentially dynamic classes $\mathcal{C}_{d}$, including vehicles, pedestrians, and cyclists, which frequently exhibit motion and require specialized handling.
 This semantic understanding enables the decomposition of the scene into two disjoint sets:
@@
 $\mathcal{G}_{i}^{s}=\{\mathcal{G}_{i}|s_{i}\in\mathcal{C}_{s}\},$
-where $\mathcal{G}_{i}^{d}$ represents dynamic Gaussians (i.e., Gaussians of potentially dynamic objects) requiring 4D modeling, and $\mathcal{G}_{i}^{s}$ represents static Gaussians.
-This semantic-guided decomposition offers several key advantages: (1) it enables specialized motion modeling for dynamic objects while preserving static scene consistency, (2) it eliminates the need for labor-intensive manual annotations, and (3) it ensures consistent motion treatment for all dynamic objects, regardless of their movement state.
+where $\mathcal{G}_{i}^{s}$ denotes the static Gaussians that require static regularization, while $\mathcal{G}_{i}^{d}$ refers to dynamic Gaussians (i.e., Gaussians associated with potentially dynamic objects) that necessitate 4D modeling.
+This semantic-guided decomposition offers several key advantages: (1) it ensures static elements remain unchanged over time while confining temporal modeling to potentially dynamic objects, (2) it enhances the reconstruction quality specifically for road, and (3) it eliminates the need for labor-intensive manual annotations.
 The decomposition lays the foundation for our two-stream optimization strategy.
-Dynamic Gaussians in $\mathcal{G}_{i}^{d}$ undergo dedicated motion refinement (Sec.
-3.3) to accurately capture movement, while static Gaussians receive geometric regularization (Sec.
-3.4) to enhance scene stability.
-Additionally, we employ an optimizable environment texture map for sky representation, which is rendered separately and composited with the Gaussian-rendered image through alpha blending, as described in [7].
-This separation allows targeted optimization without interference between dynamic and static components.
-Optimizing stationary dynamic Gaussians (e.g., parked vehicles) within a distinct subset is still more efficient than methods like DeformGS [40], which jointly model entire scene dynamics.
-This targeted approach results in a more accurate reconstruction of both dynamic and static elements.
-Figure 4: Overview of Urban4D framework.
-Given input images with semantic information, Gaussians are classified into static and potentially dynamic elements through semantic-guided decomposition.
-Dynamic objects are represented in 4D Gaussian Splatting that captures motion patterns by incorporating a learnable time embedding, with deformations predicted at desired timestamps using an MLP.
-For static Gaussians, particularly low-texture areas like ground surfaces, consistency is enforced using a KNN-based regularization mechanism.
+the static Gaussians receive geometric regularization (Sec.
+3.3) to enhance scene stability, while the potentially dynamic Gaussians in $\mathcal{G}_{i}^{d}$ undergo a dedicated motion refinement (Sec.
+3.4) to accurately capture movement.
+Additionally, we employ an optimizable environment texture map for sky representation, which is rendered separately and combined with the Gaussian-based image by alpha blending, as described in [7].
+3.3 Static Regularization
+Static invariance.
+Prior approaches (e.g., PVG) address dynamic scene reconstruction by applying timestamp-dependent transformations to each Gaussian’s 3D position $\mu$ and opacity $\alpha$.
+These transformations effectively capture dynamic motions but inevitably alter truly static parts of the scene.
+To mitigate this issue, we introduce a static consistency loss to keep these static Gaussians invariant:
+$\mathcal{L}_{\text{static}}=\sum_{i\in\mathcal{G}_{i}^{s}}w_{i}\Bigl{(}\|\mu_{i}-\mu_{i}^{\prime}\|^{2}+\|\alpha_{i}-\alpha_{i}^{\prime}\|^{2}\Bigr{)},$
+where $\mu_{i}$ and $\alpha_{i}$ denote the untransformed parts of static Gaussians, and ${\mu_{i}^{\prime}}$ and $\alpha_{i}^{\prime}$ are their transformed counterparts.
+$w_{i}$ is a semantic weight (derived from a softmax ratio) indicating the likelihood that Gaussian $i$ is truly static.
+This weighting mechanism allows fully static points ($w_{i}\approx 1$) to remain nearly unchanged, while points that are partly dynamic or uncertain ($10<w_{i}<$ 1) retain the freedom to move if necessary.
+Ground surface consistency regularization.
+In urban driving scenes, ground surfaces constitute a significant portion of the environment and typically exhibit low-texture characteristics.
+While ground-level Gaussians should theoretically share similar properties due to their homogeneous nature, enforcing strict uniformity across all ground Gaussians would be oversimplified and impractical, as real-world surfaces often contain variations and irregularities.
+The scale parameter of a Gaussian, derived from its covariance matrix, inherently encodes local geometric information analogous to surface normals [16, 9].
+A well-behaved ground surface should exhibit smooth transitions in its local geometry, making scale a particularly suitable target for regularization.
+This motivates us to regularize the scale parameters rather than other Gaussian properties.
+For each ground Gaussian $\mathcal{G}_{i}\in\mathcal{G}_{g}$ (where $\mathcal{G}_{g}\subset\mathcal{G}_{i}^{s}$ denotes the set of ground surface Gaussians), we identify its $N$ nearest neighbors:
+$\mathcal{N}_{i}=\text{KNN}(\mathcal{G}_{i},\mathcal{G}_{g},N),$ (6)
+where KNN retrieves the $N$ spatially closest Gaussians to $\mathcal{G}_{i}$ from the global set $\mathcal{G}_{g}$.
+The neighbors are determined based on the Euclidean distance between Gaussian centers $\boldsymbol{\mu}_{i}$, forming a local neighborhood for geometric consistency.
+We then introduce a local consistency loss that encourages similar scale properties within each local neighborhood:
+$\mathcal{L}_{\text{ground}}=\sum_{\mathcal{G}_{i}\in\mathcal{G}_{g}}\|\boldsymbol{s}_{i}-\frac{1}{N}\sum_{\mathcal{G}_{j}\in\mathcal{N}_{i}}\boldsymbol{s}_{j}\|_{2}^{2},$ (7)
+where $\boldsymbol{s}_{i}$ and $\boldsymbol{s}_{j}$ represents the scale parameter of the Gaussian $\mathcal{G}_{i}$ and $\mathcal{G}_{j}$.
+By regularizing the scale parameters, we effectively enforce consistency in the local surface geometry while preserving the ability to model natural surface variations.
+This approach leads to more coherent ground surface reconstruction, as similar scale parameters in a local neighborhood implicitly enforce consistent surface normal orientations, resulting in improved geometric fidelity of the ground surface representation.
 3.4 4D Gaussian Splatting Representation
-To represent our dynamic Gaussians $\{\mathcal{G}_{i}^{d}\}$ in 4D, our method extends the deformation mechanism of DeformGS [40] by introducing a learnable time embedding for each Gaussian.
-While DeformGS [40] directly maps spatial positions and time to deformations, our approach leverages temporal context through Gaussian-specific embeddings.
+While the original dynamic Gaussian approach demonstrates the capability to model dynamic objects, we aim to refine it further to better handle the remaining potentially dynamic Gaussians.
+To refine these potentially dynamic Gaussians ${\mathcal{G}_{i}^{d}}$ in 4D, our method extends the deformation mechanism of DeformGS [46] by introducing a learnable time embedding for each Gaussian.
+Unlike DeformGS [46], which directly maps spatial positions and time to deformations, our approach leverages temporal context through Gaussian-specific embeddings.
+This refinement enables a more adaptive representation of dynamic elements.
 Specifically, for each dynamic Gaussian $\mathcal{G}_{i}^{d}$, we maintain a learnable time embedding vector $\mathbf{e}_{i}\in\mathbb{R}^{D_{e}}$.
@@
 The entire refinement process is end-to-end trainable along with the main 3DGS optimization objectives.
-3.4 Ground Surface Consistency Regularization
-In urban driving scenes, ground surfaces constitute a significant portion of the environment and typically exhibit low-texture characteristics.
-While ground-level Gaussians should theoretically share similar properties due to their homogeneous nature, enforcing strict uniformity across all ground Gaussians would be oversimplified and impractical, as real-world surfaces often contain variations and irregularities.
-The scale parameter of a Gaussian, derived from its covariance matrix, inherently encodes local geometric information analogous to surface normals [12, 8].
-A well-behaved ground surface should exhibit smooth transitions in its local geometry, making scale a particularly suitable target for regularization.
-This motivates us to regularize the scale parameters rather than other Gaussian properties.
-For each ground Gaussian $\mathcal{G}_{i}\in\mathcal{G}_{g}$ (where $\mathcal{G}_{g}\subset\mathcal{G}_{i}^{s}$ denotes the set of ground surface Gaussians), we identify its $N$ nearest neighbors:
-$\mathcal{N}_{i}=\text{KNN}(\mathcal{G}_{i},\mathcal{G}_{g},N),$ (7)
-where KNN retrieves the $N$ spatially closest Gaussians to $\mathcal{G}_{i}$ from the global set $\mathcal{G}_{g}$.
-The neighbors are determined based on the Euclidean distance between Gaussian centers $\boldsymbol{\mu}_{i}$, forming a local neighborhood for geometric consistency.
-We then introduce a local consistency loss that encourages similar scale properties within each local neighborhood:
-$\mathcal{L}_{\text{ground}}=\sum_{\mathcal{G}_{i}\in\mathcal{G}_{g}}\|\boldsymbol{s}_{i}-\frac{1}{N}\sum_{\mathcal{G}_{j}\in\mathcal{N}_{i}}\boldsymbol{s}_{j}\|_{2}^{2},$ (8)
-where $\boldsymbol{s}_{i}$ and $\boldsymbol{s}_{j}$ represents the scale parameter of the Gaussian $\mathcal{G}_{i}$ and $\mathcal{G}_{j}$.
-By regularizing the scale parameters, we effectively enforce consistency in the local surface geometry while preserving the ability to model natural surface variations.
-This approach leads to more coherent ground surface reconstruction, as similar scale parameters in a local neighborhood implicitly enforce consistent surface normal orientations, resulting in improved geometric fidelity of the ground surface representation.
 3.5 Optimization Strategy
@@
 The overall loss function is formulated as follows:
-$\mathcal{L}_{\text{total}}=$ $\lambda_{1}\mathcal{L}_{\text{L1}}+\lambda_{2}\mathcal{L}_{\text{SSIM}}+\lambda_{3}\mathcal{L}_{\text{sem}}$
-$+\lambda_{4}\mathcal{L}_{\text{ground}}+\lambda_{5}\mathcal{L}_{\text{depth}}+\lambda_{6}\mathcal{L}_{\text{sky}}$ (9)
-where $\{\lambda_{i}\}_{i=1}^{6}$ are weighting coefficients balancing different loss terms. Each loss term serves a specific purpose in our optimization.
+$\mathcal{L}=$ $\lambda_{1}\mathcal{L}_{\text{L1}}+\lambda_{2}\mathcal{L}_{\text{SSIM}}+\lambda_{3}\mathcal{L}_{\text{sem}}$
+$+\lambda_{4}\mathcal{L}_{\text{static}}+\lambda_{5}\mathcal{L}_{\text{ground}}+\lambda_{6}\mathcal{L}_{\text{depth}}+\lambda_{7}\mathcal{L}_{\text{sky}},$ (11)
+where $\{\lambda_{i}\}_{i=1}^{7}$ are weighting coefficients balancing different loss terms. Each loss term serves a specific purpose in our optimization.
 Appearance Supervision. The L1 loss and SSIM loss work together to ensure accurate color reproduction and structural similarity:
@@
 where CE denotes cross-entropy loss between rendered semantic maps $S_{\text{rendered}}$ and ground-truth semantic maps $S_{\text{gt}}$.
-Ground Consistency. As detailed in Sec. 3.4, the ground consistency loss enforces local geometric similarity as defined in Eq. 8.
+Static invariance & Ground Consistency. As detailed in Sec. 3.3.
 Geometric Supervision. The inverse depth loss aligns the scene geometry with LiDAR measurements:
@@
 PVG [7] captures dynamic vehicles to some extent but suffers from noticeable blurring, particularly in the lower parts of the objects.
-In contrast, Urban4D delivers superior reconstruction quality, maintaining high fidelity and preserving clear details throughout the dynamic objects.
+In contrast, UrbanGS delivers superior reconstruction quality, maintaining high fidelity and preserving clear details throughout the dynamic objects.
 Table 1: Quantitative comparison with existing methods on the nuScenes [4] dataset.
-The best results in each group are highlighted in bold.
-Our method outperforms previous approaches across all evaluation metrics.
-Method Full Image Non-Sky Human Vehicle Dynamic
+D: DINO features, F: Optical Flow, S: Semantic map, B: Bounding box.
+* indicates methods that use ground-truth (GT) bounding boxes.
+Method Extra Inputs Full Image Non-Sky Human Vehicle Dynamic
 […]
-SUDS [30] 20.02 0.605 20.45 0.621 21.20 0.633 21.98 0.678 21.49 0.646
+SUDS [35] D+F 20.02 0.605 20.45 0.621 21.20 0.633 21.98 0.678 21.49 0.646
 DeformGS [46] 25.32 0.825 25.27 0.822 21.44 0.6456 21.72 0.700 21.62 0.684
 3DGS [17] 26.02 0.825 26.45 0.836 23.20 0.721 23.98 0.794 23.59 0.756
-EmerNeRF [38] 26.12 0.830 26.50 0.840 23.45 0.733 24.69 0.808 24.12 0.767
-PVG [7] 26.23 0.834 26.72 0.841 23.98 0.743 24.73 0.815 24.33 0.774
-Urban4D (Ours) 26.47 0.839 28.93 0.878 24.88 0.783 26.05 0.855 25.33 0.810
+EmerNeRF [44] S 26.12 0.830 26.50 0.840 23.45 0.733 24.69 0.808 24.12 0.767
+PVG [7] S 26.23 0.834 26.72 0.841 23.98 0.743 24.73 0.815 24.33 0.774
+streetGS [43] S+B 25.46 0.831 25.50 0.820 22.54 0.659 25.87 0.837 24.17 0.771
+4dgf [10] B 21.20 0.694 24.37 0.769 20.36 0.619 22.74 0.762 22.39 0.698
+OmniRe [8] S+B 26.41 0.837 26.73 0.845 23.71 0.737 25.95 0.856 25.24 0.805
+UrbanGS (Ours) S 26.92 0.848 27.61 0.861 24.92 0.767 26.01 0.838 25.43 0.818
+streetGS [43] S+$\mathrm{B}^{*}$ 25.89 0.845 26.01 0.858 22.83 0.705 26.88 0.852 25.10 0.803
+4dgf [10] $\mathrm{B}^{*}$ 27.48 0.852 27.81 0.865 25.16 0.789 27.14 0.858 26.22 0.843
+OmniRe [8] S+$\mathrm{B}^{*}$ 29.15 0.873 29.54 0.886 26.10 0.835 27.23 0.861 26.68 0.856
 4 Experiments
@@
 Following previous 3DGS-based methods to predict reliable semantic Gaussians, we use SegFormer [41] as our pre-trained segmentation model.
+Our approach builds upon PVG [7], with all parameters configured identically to its original implementation.
 The learning rate of MLP starts from $1.6\times 10^{-4}$ and decreases to $1.6\times 10^{-6}$.
-For each loss term, the weighting coefficients are empirically set to $\lambda_{1}=0.8$, $\lambda_{2}=0.2$, $\lambda_{3}=0.01$, $\lambda_{4}=0.0001$, $\lambda_{5}=0.1$ and $\lambda_{6}=0.01$.
-All experiments are conducted on a single NVIDIA V100s with 32GB memory.
+For each loss term, the weighting coefficients are empirically set to $\lambda_{1}=0.8$, $\lambda_{2}=0.2$, $\lambda_{3}=0.01$,$\lambda_{4}=0.01$, $\lambda_{5}=0.0001$, $\lambda_{6}=0.1$ and $\lambda_{7}=0.01$.
+All experiments are conducted on a single NVIDIA V100s.
 4.2 Datasets
-Our experiments are conducted on two popular autonomous driving datasets: PandaSet [7] and nuScenes [4].
-We primarily evaluate our method on nuScenes [4], which consists of 1000 scenes captured in Boston and Singapore under diverse urban scenarios with multimodal sensor data, featuring challenging weather conditions and complex urban scenes.
-We perform ablation studies on PandaSet [7], which is a comprehensive autonomous driving dataset captured in San Francisco, containing 103 sequences of synchronized LiDAR-camera data and images.
-For both datasets, we use the officially provided camera poses.
+Our experiments are conducted on two widely-used autonomous driving datasets: nuScenes [4] and PandaSet [7].
+The nuScenes dataset [4] comprises 1,000 scenes captured in Boston and Singapore under diverse urban scenarios.
+PandaSet [7] is a comprehensive dataset collected in San Francisco, containing 103 sequences with synchronized LiDAR and camera data.
+Both datasets provide ground-truth (GT) 3D bounding boxes.
+To ensure fair comparisons, we also utilize a pretrained state-of-the-art 3D multi-object tracking model, MCTrack [37], to generate predicted 3D bounding boxes for methods requiring such inputs.
+Additionally, we conducted experiments on the Waymo dataset [30]; detailed results and analyses are available in the supplementary material.
 4.3 Results and Comparisons
 Results on nuScenes [4].
-We evaluate our method against previous approaches on the nuScenes [4] dataset, with a comprehensive comparison across different scene components and object categories.
-As shown in Table 1, our method demonstrates consistent improvements over NeRF-based methods [30, 38], traditional 3DGS [13], PVG [7] and DeformGS [40] across all evaluation metrics, with significant improvements in dynamic object reconstruction.
-For full image reconstruction, our method achieves 26.47 PSNR and 0.839 SSIM, surpassing PVG [7] by 0.24 PSNR.
-The improvement is more pronounced in non-sky regions, where our approach shows a substantial gain of 2.21 PSNR (28.93 vs 26.72) and 0.037 SSIM (0.878 vs 0.841) compared to PVG [7], indicating better handling of complex geometric structures in urban environments.
-The advantages of our method are most evident in dynamic object reconstruction.
-For human instances, we achieve 24.88 PSNR and 0.783 SSIM, outperforming PVG [7] by 0.90 PSNR and 0.040 SSIM.
-In vehicle reconstruction, our method demonstrates even larger improvements, reaching 26.05 PSNR and 0.855 SSIM, surpassing PVG [7] by 1.32 PSNR and 0.040 SSIM.
-These significant gains can be attributed to our dynamic-focused semantic decomposition strategy, which enables specialized treatment of moving objects.
+We comprehensively evaluate our method against previous state-of-the-art approaches on the nuScenes [4] dataset, reporting quantitative results in Table 1.
+Our approach achieves superior performance among methods that do not rely on ground-truth (GT) bounding boxes, surpassing recent semantic-based methods (e.g., PVG [7], EmerNeRF [44]) as well as bounding-box-based approaches utilizing predicted boxes (e.g., streetGS [43], OmniRe [8], 4dgf [10]).
+Specifically, for full-image reconstruction, our method achieves 26.92 PSNR and 0.848 SSIM, significantly outperforming PVG [7] by 0.69 PSNR and 0.014 SSIM.
+In non-sky regions, our gains become even more pronounced, reaching 27.61 PSNR and 0.861 SSIM, clearly exceeding PVG by 0.89 PSNR and 0.020 SSIM.
+Such improvements highlight our method’s effectiveness in accurately reconstructing detailed urban structures.
+Remarkably, our approach not only outperforms all methods relying on predicted bounding boxes, but also surpasses the streetGS [43] method even when it leverages GT bounding boxes.
+This result demonstrates the strong robustness and effectiveness of our method, as we achieve superior accuracy without explicit bounding-box supervision.
+Additionally, methods that rely heavily on bounding boxes experience significant performance degradation when switching from GT to predicted bounding boxes; notably, the 4dgf [10] method struggles significantly under predicted bounding-box inputs, failing to converge and exhibiting very low performance.
+In dynamic object reconstruction, our method further demonstrates clear advantages.
+For human instances, we achieve 24.92 PSNR and 0.767 SSIM, substantially outperforming PVG by 0.94 PSNR and 0.024 SSIM.
+For vehicle instances, we achieve 26.01 PSNR and 0.838 SSIM, representing a significant gain of 1.28 PSNR and 0.023 SSIM compared to PVG.
+These improvements underscore our model’s strength in effectively reconstructing dynamic content in challenging urban environments, without relying on external bounding-box supervision.
 We present qualitative comparisons in Fig.
-5.
-For dynamic objects like vehicles and pedestrians, our method shows significant improvements over PVG [7] and DeformGS [40], with notably reduced motion blur and sharper object boundaries.
-This is particularly evident in regions with fast-moving objects, where previous methods tend to produce blurry reconstructions.
+1.
+For dynamic objects like vehicles and pedestrians, our method shows significant improvements over PVG [7] and DeformGS [46], with notably reduced motion blur.
 Meanwhile, our approach also demonstrates better reconstruction quality for static scene elements such as roads, preserving more detailed textures and geometric structures.
@@
 In contrast, our method achieves clearer and more consistent reconstruction across all dynamic objects.
-Table 2: Quantitative comparison with state-of-the-art methods on the PandaSet [34] dataset. We report image reconstruction and novel view synthesis metrics. The best results are shown in bold.
+Table 2: Quantitative comparison with state-of-the-art methods on the PandaSet [40] dataset. We report image reconstruction and novel view synthesis metrics. * indicates using ground-truth data
 Method Image Reconstruction Novel View Synthesis
@@
 PVG [7] 27.15 0.836 25.92 0.798
-Urban4D (Ours) 27.74 0.850 26.56 0.814
+OmniRe [8] 26.94 0.840 25.75 0.804
+UrbanGS (Ours) 28.03 0.858 26.76 0.821
+$\mathrm{OmniRe}^{*}$ [8] 29.02 0.882 27.73 0.855
 Results on PandaSet [40].
-To further validate the generalization capability of our method, we conduct additional experiments on the PandaSet [34] dataset.
-As shown in Table 2, our approach demonstrates consistently superior performance across both image reconstruction and novel view synthesis tasks.
-For image reconstruction, our method achieves 27.74 PSNR and 0.850 SSIM, outperforming PVG [7] by 0.59 PSNR and 0.014 SSIM.
-The performance advantage is even more pronounced when compared to vanilla 3D Gaussian Splatting (3DGS [13]), showing improvements of 4.07 PSNR and 0.107 SSIM.
-This significant gap highlights the effectiveness of our dynamic-focused scene decomposition strategy.
-In the more challenging novel view synthesis task, our method maintains its lead with 26.56 PSNR and 0.814 SSIM, surpassing PVG [7] by 0.64 PSNR and 0.016 SSIM.
-Notably, while all methods experience some performance degradation in novel view synthesis compared to image reconstruction, our method shows the smallest drop (1.18 PSNR vs 1.23 PSNR for PVG [7]), indicating better scene understanding and geometry modeling.
-The consistent performance improvements across different datasets and evaluation metrics demonstrate the robustness and generalizability of our approach across varied urban scenes and capture conditions.
+We further evaluate our method on the PandaSet dataset [40], comparing it against recent state-of-the-art methods in both image reconstruction and novel view synthesis tasks (Table 2).
+Our approach consistently outperforms previous methods that rely on predicted bounding boxes or do not use bounding boxes at all, including OmniRe [8], PVG [7], EmerNeRF [44], and 3DGS [17].
+Although methods utilizing ground-truth bounding boxes achieve higher performance, our method demonstrates superior robustness and effectiveness under realistic conditions (i.e., without ground-truth bounding boxes).
+These results highlight the effectiveness of our proposed approach in accurately reconstructing urban scenes.
+For additional qualitative results, please refer to the Supplementary.
 4.4 Ablation Study
-Effectiveness of consistency regularization.
-To validate the effectiveness of our ground surface consistency module, we conduct extensive experiments on the PandaSet [34] dataset.
-Table 3 shows the quantitative results of different variants of our method.
-The baseline model without road consistency achieves 26.31 PSNR and 0.810 SSIM.
-After incorporating our consistency constraint, the performance significantly improves by 0.25 dB in PSNR (26.56 vs.
-26.31) and 0.004 in SSIM (0.814 vs.
-0.810).
-This demonstrates that enforcing local geometric and appearance consistency on ground surfaces effectively enhances the reconstruction quality.
-We further investigate the impact of different numbers of nearest neighbors in our module.
-As shown in Table 3, using 16 neighbors yields the best performance (26.56 PSNR), outperforming both smaller (k=8, 26.45 PSNR) and larger (k=32, 26.48 PSNR) neighbor sizes.
-We hypothesize that 8 neighbors are insufficient to capture the local surface structure, while 32 neighbors may introduce noise from points that are too far away.
-The optimal choice of 16 neighbors provides a good local structure preservation.
-Table 3: Ablation study on road consistency module. We evaluate different design choices and parameter settings of our road consistency constraint on PandaSet [7].
+We conduct an ablation study to evaluate the contribution of each module in our method on the PandaSet [7]. The results are summarized in Table 3.
+Effect of Static Invariance.
+The absence of the static invariance module results in a notable decline in performance, with PSNR and SSIM dropping to 26.56 and 0.814, respectively.
+This underscores the importance of incorporating static priors to improve the reconstruction of static regions, thereby ensuring more robust and accurate results.
+Effect of Road Consistency.
+Removing the road consistency module causes a performance drop.
+This demonstrates that enforcing road consistency is essential for preserving the geometric and textural integrity of roads, which significantly enhances the overall reconstruction quality.
+Effect of 4D Representation.
+Excluding the 4D representation leads to the most significant performance degradation, with PSNR reducing to 26.43 and SSIM dropping to 0.810.
+This highlights the critical contribution of the 4D representation in modeling potentially dynamic objects and handling temporal variations, which are essential for reconstructing complex urban scenes.
+Visualization of Static Regularization.
+Figure 5 presents a visual comparison of the effect of static regularization on Waymo [30].
+When the Static Invariance and Road Consistency modules are removed, the reconstructed ground becomes significantly more blurred.
+In contrast, incorporating these modules results in much clearer and more visually pleasing reconstructions, demonstrating their effectiveness.
+Table 3: Ablation study on each module. We evaluate the effect of each module on PandaSet [7].
 […]
-Baseline (w/o road consistency) 26.31 0.810
-KNN-8 26.45 0.812
-KNN-16 (Ours) 26.56 0.814
-KNN-32 26.48 0.812
-Table 4: Ablation study on time embedding. We report PSNR and SSIM on the PandaSet [34].
-[…]
-w/o Time Embeddings 26.40 0.810
-w/ Time Embeddings 26.56 0.814
-Effectiveness of Time Embeddings.
-We also validate the effectiveness of our temporal modeling strategy through ablation studies.
-As shown in Table 4, equipping each Gaussian with a learnable time embedding improves the rendering quality (PSNR: 26.40 → 26.56, SSIM: 0.810 → 0.814) on the PandaSet [34] dataset.
-This demonstrates that point-specific temporal context helps capture more accurate deformation patterns compared to directly mapping spatial positions to deformations.
+Baseline (w/o static invariance) 26.56 0.814
+Baseline (w/o road consistency) 26.60 0.816
+Baseline (w/o 4D Representation) 26.43 0.810
+UrbanGS(Ours) 26.76 0.821
+Figure 5: Ablation study on the effectiveness of static regularization. Results without static regularization (left) are blurry, while adding it (right) produces sharper details.
 5 Conclusions
-In conclusion, Urban4D provides a novel semantic-guided decomposition and 4D representation strategy for reconstructing dynamic urban scenes.
-By leveraging 2D semantic information, our approach effectively separates static and potentially dynamic components.
-We introduced specialized processing methods for different elements: a 4D Gaussian Splatting representation for dynamic objects and KNN-based consistency regularization for the ground surface.
-Both quality and quantitative results demonstrated that Urban4D improves rendering quality across various scene components, particularly for dynamic objects and low-texture regions.
+In conclusion, UrbanGS provides a novel semantic-guided decomposition strategy for reconstructing urban scenes.
+By leveraging 2D semantic information, our approach effectively separates static elements from potentially dynamic components.
+We introduced specialized processing methods for different elements: A static Invariance module to leverage static priors to enhance the reconstruction of stationary regions and KNN-based consistency regularization for the ground surface.
+And a 4D Gaussian Splatting representation for potentially dynamic objects.
+Both qualitative and quantitative results demonstrated that UrbanGS improves rendering quality across various scene components.